OpenStudy (anonymous):

Write 3 - 5 + 7 - 9 + 11 - 13 + 15 in sigma notation.

Parth (parthkohli):

\[\textbf{HINT}\]You'd have to use the powers of \(-1\) because there are sign changes in your series. Also, use the property of odd numbers.

OpenStudy (anonymous):

Meaning....? Sorry I'm really lost here

Parth (parthkohli):

You know that every odd number is in the form \(2n+1\) where \(n\) is some integer.

OpenStudy (anonymous):

so the bottom would be n-1?

OpenStudy (anonymous):

\[\sum_{n-1}^{?}\] right?

OpenStudy (skullpatrol):

No.

Parth (parthkohli):

I'm not really sure how I'd explain this to you, so here's two exemplary hints.\[\sum_{\Large{n =0}}^{\Large 5}(-1)^n = (-1)^0 + (-1)^1 + (-1)^2\cdots(-1)^5 = 1 +(-1) + 1 \cdots(-1)\]and\[\sum_{\Large{n = 1}}^{\Large 5} 2n + 1 = 2(1) + 1 + 2(2) + 1\cdots2(5) + 1 = 3 + 5 + \cdots+11\]

Parth (parthkohli):

Another (and probably the last) hint: \((-1)^1\times 5 = -5\) and \((-1)^2 \times 5 = 5\).

Parth (parthkohli):

Now can you think about it a little? I'm ready to support you through the rest.

OpenStudy (anonymous):

I just found where we weren't suppose to do that problem. Sorry all!

Parth (parthkohli):

lol okay

Parth (parthkohli):

You're welcome BTW. =)

OpenStudy (anonymous):

Thanks for trying tho!